write a program to check if a student is eligible for admission in professional course based on the following criteria:


Sum of marks in any two subjects >= 100 and M+P+C>=180


1
Expert's answer
2021-12-13T00:03:23-0500
M=int(input("Enter marks for Maths: "))
P=int(input("Enter marks for Physics: "))
C=int(input("Enter marks for Chemistry: "))


if(((M+P>=100) or (M+C>=100) or (P+C>=100)) and (M+P+C>=180)):
    print("TRUE")
else:
    print("FALSE")